# Efficient absorbing boundary conditions for conservation laws: Application to the piston problem of gas dynamics

### v1

@@ Abstract
In this work, we address the imposition of outflow boundary conditions for one-dimensional conservation laws. While a highly accurate numerical solution can be obtained in the interior of the domain, boundary discretization can lead to unphysical reflections. We investigate and implement some classes of relaxation methods and far-field operators to deal with this problem without significantly increasing the size of the computational domain. Relations are established between these techniques, and extensions of them are explored. In particular, we introduce a simple and robust relaxation method with a matrix-valued weight function that selectively absorbs outgoing waves. As a challenging model problem, we consider the Lagrangian formulation of the Euler equations for an isotropic gas with inflow boundary conditions determined by an oscillating piston.
@@ 1 Introduction
Hyperbolic systems of partial differential equations (PDEs) are characterized by finite propagation speed of information in their solutions. At a fixed time, and with given bounded and compactly supported initial data, we expect the relevant information of a phenomenon described by a system of conservation laws to be concentrated inside a bounded subset of the domain. Nevertheless, when the physical domain is infinite, increasingly extensive computational domains might be necessary, which can become prohibitively expensive. Even though this has been partially alleviated by the fast development of computer hardware, to accurately capture the long-time behavior of solutions still poses a significant challenge. As a consequence, a variety of absorbing boundary conditions (ABCs) have been proposed to tackle this problem (see, e.g., [8, 13] and references therein).
@@
In this manuscript, we focus on the so-called relaxation and far-field operators methods, briefly outlined below. These are families of ABCs that could, in principle, be applied to any system of hyperbolic PDEs. Both of them rely on the use of a region of the domain where the numerical solution is damped or forced towards a given function. To avoid ambiguity, we will refer to this region as the sponge layer throughout this work (see Figure 1).
@@
Larsen and Dancy [25] introduced the relaxation method to apply ABCs to their discretization of Peregrine’s Boussinesq type equations. The method was first presented as a postprocessing technique where the numerical solution is divided by an increasing set of positive numbers inside the sponge layer (often referred to as relaxation zone in the literature) at each step of the time integration procedure. A Fourier mode-based reflection analysis of a linearization of Peregrine’s equations suggested that the absorbing properties of this technique improve drastically as the ratio between the the length of the sponge layer and the wavelength increases. Due to the versatility and efficiency of this method, it has been widely used for both absorption and generation of waves in water wave modeling, where validation tests often consist of solitary waves with a known characteristic length that can be contained inside the sponge layer [28, 6, 17]. Furthermore, several works that formalize this technique and analyze its non-reflecting properties have been proposed, e.g. [30, 4, 5, 14].
@@
Another popular approach for the development of ABCs is the modification of the system of governing equations, such that damping of outgoing waves is a characteristic behavior of the analytic solution. In their seminal work [16], Israeli and Orszag discuss the addition of reaction and diffusion terms with variable strength inside the sponge layer. They further explore the idea of absorbing waves in just one direction for the wave equation by modifying its second-order scalar and first-order system forms. However, it was later shown by Karni [19] that such modifications unavoidably generate reflections that propagate into the computational domain. Instead, Karni proposes the use of far-field slowing-down and damping operators that, by preserving the eigenstructure of the flux Jacobian, absorb waves in a prescribed direction for systems of hyperbolic PDEs. A generalization of Karni’s slowing-down and damping operators for linear systems was introduced by Appelö and Colonius [1], where high-order viscosity terms are added in the sponge layer.
@@
As a benchmarking application for the ABCs discussed in this work, we consider the one-dimensional Euler equations in their Lagrangian formulation. These equations are used to model the piston problem of gas dynamics, where the solution is known to consist of quasi-periodic nonlinear plane waves with strength proportional to the amplitude of the source [40]. In particular, we study the strongly nonlinear case, where shock formation occurs in the near-field and recurrent interaction between the first and second shocks of the wave train takes place in the far-field, as demonstrated by Inoue and Yano [15]. A central scheme discretization for this problem and its multi-fluid counterpart in Lagrangian coordinates was put forward by Fazio and Russo [7], where the prescription of second-order accurate inflow boundary conditions at the piston, used herein, is described.
@@
  • a simple relation between the relaxation method and far-field damping operators ABCs for general hyperbolic systems of conservation laws (see Sections 3.1 and 3.3);
  • A generalization of the relaxation method to improve its absorbing properties without significantly increasing its computational cost (see Section 3.2);
@@
  • a detailed performance comparison of different state-of-the-art ABCs for the Lagrangian formulation of the Euler equations with emphasis on the piston problem of gas dynamics (see Section 5).
This manuscript is organized as follows. In Section 2 some existing sponge layer-based ABCs are reviewed. In Section 3, we introduce extensions of the techniques discussed in Section 2. In Section 4 we recall the compressible Euler equations for an isotropic gas, their Lagrangian formulation, and the piston problem of gas dynamics. In Section 5 we assess the performance of the different ABCs in the context of the piston problem.
The code used for the development of this work is available in the reproducibility repository [31].
@@ 2 Sponge layer-based ABCs
Here $\mathbf{q}\in L^{1}([0,T]\times[0,+\infty[,^{N})$, and the Jacobian matrix of $\mathbf{f}$ can be diagonalized as $\mathbf{f}^{\prime}(\mathbf{q})=\mathbf{R}\mathbf{\Lambda}{\mathbf{R}}^{-1}$, where the columns of $\mathbf{R}$ are the right eigenvectors $\{\mathbf{r}_{i}(\mathbf{q})\}_{i=1}^{N}$ of $\mathbf{f}^{\prime}(\mathbf{q})$, and $\boldsymbol{\Lambda}$ is a diagonal matrix containing its real eigenvalues $\{\lambda_{i}(\mathbf{q})\}_{i=1}^{N}$. The linear case, where $\mathbf{f}(\mathbf{q})=\mathbf{A}\mathbf{q}$ for some constant matrix $\mathbf{A}\in^{N\times N}$,
@@
The computational domain and sponge layer will be the intervals $[0,x_{s}[$ and $[x_{s},x_{\infty}]$ respectively, and we suppose there is a far-field steady-state $\overline{\mathbf{q}}\in^{N}$ such that $\mathbf{q}(0,x)=\overline{\mathbf{q}}$ for $x>x_{s}$. Our goal is to impose ABCs in the sponge layer so that perturbations of $\overline{\mathbf{q}}$ in the solution are absorbed before they reach the right boundary $x_{\infty}$, as depicted in Figure 1. Regarding the left boundary, we assume the existence of a function $\mathbf{q}_{BC}(\cdot)$ such that well-posedness is guaranteed with the inflow boundary condition $\mathbf{q}(t,0)=\mathbf{q}_{BC}(t)$.
@@ 2.1 Slowing-down and damping far-field operators (SDO)
In the work [19], Karni proposes to impose ABCs for the linear system (2) by solving the modified system
@@
instead. $\mathbf{S}$ and $\mathbf{D}$ are referred to as slowing-down and damping far-field operators respectively, and are given by
$\mathbf{S}(x)=\mathbf{R}{\bf\Lambda_{S}}(x)\mathbf{R}^{-1},\quad\mathbf{D}(x)=\mathbf{R}{\bf\Lambda_{D}}(x)\mathbf{R}^{-1},$ (4)
where ${\bf\Lambda_{S}}(x)$ and ${\bf\Lambda_{D}}(x)$ are diagonal matrices with entries $\{s_{i}(x)\lambda_{i}\}_{i=1}^{N}$ and $\{d_{i}(x)s_{i}(x)\lambda_{i}\}_{i=1}^{N}$.
After writing (3) in terms of the characteristic variables $\mathbf{w}=\mathbf{R}^{-1}\mathbf{q}$, and if $s_{i}$ and $d_{i}$ are constant, we have $N$ decoupled advection-reaction equations
$\partial_{t}w_{i}+s_{i}\lambda_{i}\partial_{x}w_{i}=-d_{i}s_{i}\lambda_{i}(w_{i}-\overline{w}_{i}),$ (5)
with explicit solution
$w_{i}(x,t)=\exp(-d_{i}x)\left[\overset{\circ}{w}_{i}(x-s_{i}\lambda_{i}t)-\overline{w}_{i}\right],$ (6)
where $\overset{\circ}{\mathbf{q}}(x)$ is the initial condition for the homogeneous problem. Thus, the speed of propagation and strength of waves corresponding to the $i$-th characteristic field will be reduced at the rates $s_{i}$ and $d_{i}$ respectively. The rationale behind these modifications is that outgoing waves (corresponding to $\lambda_{j}>0$ for an outlet at the right of the domain) can be both slowed down and damped by choosing $s_{j}<1$ and $d_{j}>0$ inside the sponge layer. This way, outgoing waves neither reach the boundary nor reflect. Furthermore, physical waves that are expected to propagate back into the computational domain remain unperturbed.
Karni shows that by choosing the operators as in Equation (4) there are no reflections at an interface point $x$ such that $s_{j}({x}^{-})\neq s_{j}({x}^{+})$ or $d_{j}({x}^{-})\neq d_{j}({x}^{+})$ at the analytical level. However, at the numerical level, reflections proportional to the jump in the coefficients $s_{j}$ and $d_{j}$ are expected [19]. Therefore, to absorb the $j$-th wave, $s_{j}(x):=\Gamma(x)$ and $d_{j}(x):=d(x)$ should be smooth functions as depicted in Figure 3. Moreover, coarser grids and narrower sponge layers will lead to larger unwanted reflections. Since we are interested in absorbing right-going waves, we can set $s_{j}\equiv 1$, $d_{j}\equiv 0$ for $\lambda_{j}\leq 0$, and $0\leq s_{j}(x)\leq 1$, $d_{j}(x)\geq 0$ for $\lambda_{j}>0$. In the following, we will use the abbreviation SDO to refer to this particular choice.
If all the waves are slowed down and damped at the same rate instead, System (3) becomes
$\partial_{t}\mathbf{q}+s(x)\mathbf{A}\partial_{x}\mathbf{q}=-d(x)s(x)\mathbf{A}\left(\mathbf{q}-\overline{\mathbf{q}}\right).$ (7)
@@
Damping source terms with a similar form have been applied previously to, e.g., the wave equation [16], the linearized Euler equations [9], and the Schrödinger equation [23]. For the sake of comparison, we will also consider (7) and refer to it as scalar slowing-down and damping operators (S-SDO).
@@
Even though the SDO technique was introduced for a linear system, a natural way to use it for nonlinear systems is to compute the far-field terms (4) (now dependent on $\mathbf{q}$) locally [18]. This will be the approach considered in Section 5 for the nonlinear equations (39). Additionally, an alternative extension to the nonlinear setting is proposed in Section 3.4.
@@ 2.2 The relaxation method (RM)
The relaxation method is a postprocessing technique where the numerical solution is forced towards a prescribed function after each time step. The process is described in Algorithm 1, where the so-called weight function $\Gamma(x)$ decreases from $1$ to 00 in the sponge layer, as depicted in Figure 3. In the relaxation step, $\mathbf{q}$ is replaced by a convex combination of the numerical solution $\mathbf{q}^{*}$ and some target function, which in this case is the constant far-field state $\overline{\mathbf{q}}$.
@@
$\mathbf{q}^{*}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with initial data $\mathbf{q}^{[n]}$triangleright Hyperbolic step
$\mathbf{q}^{[n+1]}\leftarrow\Gamma(x)\mathbf{q}^{*}+(1-\Gamma(x))\overline{\mathbf{q}}$triangleright Relaxation step
@@
It is well known that the performance of this ABC depends on the choice of the weight function [3]. By considering the method as a nodal filtering procedure, Engsig-Karup [5] shows that a necessary condition to preserve the continuity of the modal coefficients through elements is that
@@
which is similar to the smooth transition requirement on the slowing-down and damping functions proposed by Karni [19]. Based on this observation, in [4] the weight function
@@
In [30], Mayer et al. propose the weight function
@@
The accuracy of this method also depends on the sponge layer length. A customary rule of thumb to design sponge layers for the relaxation method is the requirement $|x_{\infty}-x_{s}|>2L$, where $L$ is the largest wavelength present in the solution [30, 4, 3]. While this restriction might be inconvenient for problems involving long waves, it is not unique to the relaxation method. For instance, Perić and Abdel-Maksoud [34] propose the same requirement on the sponge layer for their linear and quadratic damping far-field operators.
In order to improve the properties of this method, a popular approach is to define a specialized weight function. Hu et al. [14] choose the weight function at each time step such that the mass absorbed artificially is equal to the mass entering the sponge layer. They report that, with this approach, acceptable results are attained with the looser restriction $|x_{\infty}-x_{s}|>0.4L$ for the 2D Navier-Stokes equations with free surface. In [36], Seng notes that the damping effect of the relaxation method increases dramatically as the time step size is refined. To mitigate this phenomenon, the author proposes a weight function that is updated dynamically depending on the distance between the numerical and target solutions.
A variation of the relaxation method was proposed by Wasistho et al. [39] for the 2D Navier-Stokes equations, where the authors proposed the application of relaxation after each stage of the time integration with a Runge-Kutta scheme. Some numerical results concerning this approach are presented in Section 5.3.

### v2

# A Unified Framework for Sponge-Layer Relaxation Methods and Damping Operators for Conservation Laws: Application to the Piston Problem of Gas Dynamics

@@ Abstract
This work addresses the imposition of outflow boundary conditions for one-dimensional conservation laws. While a highly accurate numerical solution can be obtained in the interior of the domain, boundary discretization can lead to unphysical reflections. We investigate and implement some classes of relaxation methods and far-field operators to deal with this problem without significantly increasing the size of the computational domain. We formulate these methods within a framework that allows to reveal relationships among them, and to propose novel extensions. In particular, we introduce a simple and robust relaxation method with a matrix-valued weight function that selectively absorbs outgoing waves. As a challenging model problem, we consider the Lagrangian formulation of the Euler equations for a polytropic gas with inflow boundary conditions determined by an oscillating piston.
keywords:
Absorbing boundary conditions, Hyperbolic PDEs, Piston problem, Finite volume method
@@ 1 Introduction
Hyperbolic systems of partial differential equations (PDEs) are characterized by finite propagation speed of information in their solutions. At a fixed time, and with given bounded and compactly supported initial data, we expect the relevant information of a phenomenon described by a system of conservation laws to be concentrated inside a bounded subset of the domain. Nevertheless, when the physical domain is infinite, increasingly extensive computational domains might be necessary in numerical simulations, which can become prohibitively expensive. Even though this has been partially alleviated by the fast development of computer hardware, to accurately capture the long-time behavior of solutions still poses a significant challenge. As a consequence, a variety of techniques have been proposed to reduce spurious reflections at artificial boundaries (see, e.g., [giles_non-reflecting_nodate, hu_absorbing_2004] and references therein).
@@
In this manuscript, we focus on the so-called relaxation and far-field operators methods, briefly outlined below. These are families of absorbing boundary conditions (ABCs)  [^1] Although the term “absorbing boundary condition” (ABC) is traditionally reserved for conditions imposed directly at the boundary, we adopt the broader engineering usage and refer to the zonal approaches considered in this work as ABCs. that could, in principle, be applied to any system of hyperbolic PDEs. Both of them rely on the use of a region of the domain where the numerical solution is damped or forced towards a given function. To avoid ambiguity, we will refer to this region as the sponge layer throughout this work (see Figure 1).
@@
Larsen and Dancy [larsen_open_1983] introduced the relaxation method to apply ABCs to their discretization of Peregrine’s Boussinesq type equations. The method was first presented as a postprocessing technique where the numerical solution is divided by an increasing set of positive numbers inside the sponge layer (often referred to as relaxation zone in the literature) at each step of the time integration procedure. A Fourier mode-based reflection analysis of a linearization of Peregrine’s equations suggested that the absorbing properties of this technique improve drastically as the ratio between the length of the sponge layer and the wavelength increases. Due to the versatility and efficiency of this method, it has been widely used for both absorption and generation of waves in water wave modeling, where validation tests often consist of solitary waves with a known characteristic length that can be contained inside the sponge layer [madsen_boussinesq-type_2003, escalante_efficient_2019, jacobsen_wave_2012]. Furthermore, several works that formalize this technique and analyze its non-reflecting properties have been proposed, e.g. [mayer_fractional_1998, engsig-karup_nodal_2006, engsig-karup_unstructured_nodate, hu_numerical_2015].
@@
Another popular approach for the development of ABCs is the modification of the system of governing equations, such that damping of outgoing waves is a characteristic behavior of the analytic solution. In their seminal work [israeli_approximation_1981], Israeli and Orszag discuss the addition of reaction and diffusion terms with variable strength inside the sponge layer. They further explore the idea of absorbing waves in just one direction for the wave equation by modifying its second-order scalar and first-order system forms. However, it was later shown by Karni [karni_far-field_1996] that such modifications unavoidably generate reflections that propagate into the computational domain. Instead, Karni proposes the use of far-field slowing-down and damping operators that, by preserving the eigenstructure of the flux Jacobian, absorb waves in a prescribed direction for systems of hyperbolic PDEs. A generalization of Karni’s slowing-down and damping operators for linear systems was introduced by Appelö and Colonius [appelo_high-order_2009], where high-order viscosity terms are added in the sponge layer.
@@
As a benchmarking application for the ABCs discussed in this work, we consider the one-dimensional Euler equations in their Lagrangian formulation. These equations are used to model the piston problem of gas dynamics, where the solution is known to consist of quasi-periodic nonlinear plane waves with strength proportional to the amplitude of the source [webster_finiteamplitude_1977]. In particular, we study the strongly nonlinear case, where shock formation occurs in the near-field and recurrent interaction between the first and second shocks of the wave train takes place in the far-field, as demonstrated by Inoue and Yano [inoue_propagation_1993]. A central scheme discretization for this problem and its multi-fluid counterpart in Lagrangian coordinates was put forward by Fazio and Russo [fazio_central_2010], where the prescription of second-order accurate inflow boundary conditions at the piston, used herein, is described.
@@
  • the formulation of a unified framework connecting relaxation methods and far-field (damping-operator) approaches for general hyperbolic systems of conservation laws (see Sections 3.1 and 3.3);
  • a generalization of the relaxation method to improve its absorbing properties without compromising robustness or significantly increasing its computational cost (see Section 3.2);
@@
  • a quantitative comparison of wave reflection for several state-of-the-art ABCs applied to the Lagrangian formulation of the Euler equations, with emphasis on the piston problem of gas dynamics (see Section 5);
This manuscript is organized as follows: Section 2 reviews existing sponge layer-based ABCs. Section 3 introduces extensions of these techniques, including improved relaxation methods and new damping operators. Section 4 recalls the compressible Euler equations in Lagrangian form and the piston problem. Section 5 provides a performance comparison of the different ABCs.
The code used for the development of this work is available in the reproducibility repository [Munoz_repository_2024].
@@ 2 Sponge layer-based ABCs
Here $\mathbf{q}\in L^{1}([0,T]\times[0,+\infty[,\mathbb{R}^{N})$, and the Jacobian matrix of $\mathbf{f}$ can be diagonalized as $\mathbf{f}^{\prime}(\mathbf{q})=\mathbf{R}\bm{\Lambda}{\mathbf{R}}^{-1}$, where the columns of $\mathbf{R}$ are the right eigenvectors $\{\mathbf{r}_{i}(\mathbf{q})\}_{i=1}^{N}$ of $\mathbf{f}^{\prime}(\mathbf{q})$, and $\bm{\Lambda}$ is a diagonal matrix containing its real eigenvalues $\{\lambda_{i}(\mathbf{q})\}_{i=1}^{N}$. The linear case, where $\mathbf{f}(\mathbf{q})=\mathbf{A}\mathbf{q}$ for some constant matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$,
@@
The computational domain and sponge layer will be the intervals $0,x_{s}[$ and $x_{s},x_{\infty}]$ respectively, and we suppose there is a far-field steady-state $\overline{\mathbf{q}}\in\mathbb{R}^{N}$ such that $\mathbf{q}(0,x)=\overline{\mathbf{q}}$ for $x>x_{s}$. Our goal is to impose ABCs in the sponge layer so that perturbations of $\overline{\mathbf{q}}$ in the solution are absorbed before they reach the right boundary $x_{\infty}$, as depicted in Figure 1. Regarding the left boundary, we assume the existence of a function $\mathbf{q}_{BC}(\cdot)$ such that well-posedness is guaranteed with the inflow boundary condition $\mathbf{q}(t,0)=\mathbf{q}_{BC}(t)$.
@@ 2.1 Slowing-down and damping far-field operators (SDO)
In the work [karni_far-field_1996], Karni proposes to solve the linear system
@@
instead of the original system (2) to impose ABCs. Here, $\mathbf{S}(x)$ and $\mathbf{D}(x)$ are referred to as slowing-down and damping far-field operators respectively, and are given by
$\mathbf{S}(x)=\mathbf{R}{\bm{\Lambda}_{\bf S}}(x)\mathbf{R}^{-1},\quad\mathbf{D}(x)=\mathbf{R}{\bm{\Lambda}_{\bf D}}(x)\mathbf{R}^{-1},$ (4)
where ${\bm{\Lambda}_{\bf S}}(x)$ and ${\bm{\Lambda}_{\bf D}}(x)$ are diagonal matrices with entries $\{s_{i}(x)\lambda_{i}\}_{i=1}^{N}$ and $\{d_{i}(x)s_{i}(x)|\lambda_{i}|\}_{i=1}^{N}$. The idea behind this approach is to define the operators $\mathbf{S}$ and $\mathbf{D}$ such that the solutions of (3) coincide with those of (2) in the computational domain, while outgoing waves are slowed down and damped in the sponge layer. In [karni_far-field_1996], this is achieved by assuming each $d_{i}$ ($s_{i}$) to be a piecewise constant function that is 0 (11) in the computational domain and increases (decreases) to a prescribed value in the sponge layer. Herein, we motivate this approach at the continuous level, which will be of further use in Section A.
Let $\Gamma_{d},\Gamma_{s}\in\mathcal{C(\mathbb{R})}$ be two smooth functions that strictly decrease from $1$ to 0 in the sponge layer, and are constant everywhere else. Next, for each $i\in\{1,\ldots,N\}$, we define
$d_{i}(x)=\sigma_{i}(1-\Gamma_{d}(x)),\quad s_{i}(x)=\delta_{i}\Gamma_{s}(x),$ (5)
where $\sigma_{i}\geq 0$ is a dimensional parameter (with units of inverse time) that indicates the maximum damping rate that will be attained in the sponge layer, and $\delta_{i}\in\{0,1\}$ is an indicator variable that allows to switch off the slowing-down effect for the $i$-th characteristic field if desired. By writing System (3) in terms of the characteristic variables $\mathbf{w}=\mathbf{R}^{-1}\mathbf{q}$, we get $N$ initial value problems of the form
[…]
Let us track the solution of (6) accross the characteristic curve $X(t)$ such that
$X^{\prime}(t)=\lambda_{i}s_{i}(X(t)),\quad X(0)=x_{0}.$ (7)
This gives us the separable ODE
$\frac{d}{dt}w_{i}(t,X(t))=-|\lambda_{i}|d_{i}(X(t))s_{i}(X(t))w_{i}(t,X(t)),$ (8)
which can be solved explicitly to yield
[…]
For the sake of clarity, let us assume that $s_{i}(x)\equiv 1$ (see Remark 1). Then, we get an explicit expression for the characteristic variable $w_{i}$
[…]
which, taking the change of variables $z=x-\lambda_{i}(t-r)$, becomes
[…]
Consequently, as long as $x<x_{s}$, the information of each characteristic field is transported without modification and it decays exponentially at a non-local rate inside the sponge layer.
In [karni_far-field_1996], it is proposed to leave the left-going waves unchanged, i.e., to set
$d_{i}(x)=0,\;s_{i}(x)=1,\quad$ $\text{if }\lambda_{i}<0.$ (12)
This way, physical waves that are expected to propagate back into the computational domain remain unperturbed. In the linear case, if no information is in the sponge layer at the initial time, there is nothing to be propagated back into the computational domain through the left-going waves, so the choice of $d_{i}$ and $s_{i}$ for $\lambda_{i}<0$ does not affect the solution. In the nonlinear case, however, different characteristic fields might interact with each other in the sponge layer, and the choice (12) could allow to preserve physical incoming information, as will be demonstrated in Section 5.3. In the following, we will use the abbreviation SDO to refer to the modified system (3) with (5), and (12).
Alternatively, one could choose to damp not only the outgoing waves, but all the characteristic fields inside the sponge layer. If given $d(x)$ and $s(x)$ defined as in (5), we take $d_{i}(x)=\frac{d(x)}{|\lambda_{i}|}$ and $s_{i}(x)=s(x)$ for every $i\in\{1,\ldots,N\}$, from (3) we recover the familiar system
$\partial_{t}\mathbf{q}+s(x)\mathbf{A}\partial_{x}\mathbf{q}$ $=-\mathbf{R}\operatorname{diag}\left(d(x)s(x)\right)\mathbf{R}^{-1}\left(\mathbf{q}-\overline{\mathbf{q}}\right)$ (13a)
$=-d(x)s(x)\left(\mathbf{q}-\overline{\mathbf{q}}\right).$ (13b)
@@
Damping source terms with a similar form have been applied previously to, e.g., the wave equation [israeli_approximation_1981], the linearized Euler equations [goodrich_comparison_2008], and the Schrödinger equation [kosloff_absorbing_1986]. For the sake of comparison, we will also consider (13) and refer to it as scalar slowing-down and damping operators (S-SDO).
@@
Even though the SDO technique was introduced for a linear system, a natural way to use it for nonlinear systems is to compute the far-field terms (4) (now dependent on $\mathbf{q}$) locally [karni_accelerated_1992]. This will be the approach considered in Section 5 for the nonlinear equations (45). Additionally, an alternative extension to the nonlinear setting is proposed in Section 3.4.
Remark 1.
In the general case where $s_{i}$ is not constant, at any given time $t$ and position $x$, it is possible to track the characterstic curves $x=X(t)$ back to $x_{0}$. It suffices to solve (7), which yields $t=\int_{x_{0}}^{x}\frac{1}{\lambda_{i}s_{i}(z)}dz$. Then, by defining $T(x)=\int_{-\infty}^{x}\frac{1}{\lambda_{i}s_{i}(z)}dz$ and noting that $T$ is invertible in $0,x_{\infty}[$, we can write $x=T^{-1}(t+T(x_{0}))$ and $x_{0}=T^{-1}(T(x)-t)$.
@@ 2.2 The relaxation method (RM)
The relaxation method is a postprocessing technique where the numerical solution is forced towards a prescribed function after each time step. The process is described in Algorithm 1, where the so-called weight function $\Gamma(x)$ decreases from $1$ to 0 in the sponge layer, as depicted in Figure 4. In the relaxation step, $\mathbf{q}$ is replaced by a convex combination of the numerical solution $\mathbf{q}^{*}$ and some target function, which in this case is the constant far-field state $\overline{\mathbf{q}}$.
@@
$\mathbf{q}^{*}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with
 initial data $\mathbf{q}^{[n]}$ $\triangleright$ Hyperbolic step
$\mathbf{q}^{[n+1]}\leftarrow\Gamma(x)\mathbf{q}^{*}+(1-\Gamma(x))\overline{\mathbf{q}}$ $\triangleright$ Relaxation step
@@
It is well known that the performance of this ABC depends on the choice of the weight function [choi_performance_2020]. By considering the method as a nodal filtering procedure, Engsig-Karup [engsig-karup_unstructured_nodate] shows that a necessary condition to preserve the continuity of the modal coefficients through elements is that
@@
which is similar to the smooth transition requirement on the slowing-down and damping functions proposed by Karni [karni_far-field_1996]. Based on this observation, in [engsig-karup_nodal_2006] the weight function
@@
In [mayer_fractional_1998], Mayer et al. propose the weight function
@@
The accuracy of this method also depends on the sponge layer length. A customary rule of thumb to design sponge layers for the relaxation method is the requirement $|x_{\infty}-x_{s}|>2L$, where $L$ is the largest wavelength present in the solution [mayer_fractional_1998, engsig-karup_nodal_2006, choi_performance_2020]. While this restriction might be inconvenient for problems involving long waves, it is not unique to the relaxation method. For instance, Perić and Abdel-Maksoud [peric_reliable_2016] propose the same requirement on the sponge layer for their linear and quadratic damping far-field operators.
In order to improve the properties of this method, a popular approach is to define a specialized weight function. Hu et al. [hu_numerical_2015] choose the weight function at each time step such that the mass absorbed artificially is equal to the mass entering the sponge layer. They report that, with this approach, acceptable results are attained with the looser restriction $|x_{\infty}-x_{s}|>0.4L$ for the 2D Navier-Stokes equations with free surface. In [seng_slamming_2012], Seng notes that the damping effect of the relaxation method increases dramatically as the time step size is refined. To mitigate this phenomenon, the author proposes a weight function that is updated dynamically depending on the distance between the numerical and target solutions.
A variation of the relaxation method was proposed by Wasistho et al. [wasistho_simulation_1997] for the 2D Navier-Stokes equations, where the authors proposed the application of relaxation after each stage of the time integration with a Runge-Kutta scheme. Some numerical results concerning this approach are presented in Section 5.3.